# A Self-Improvable Polymer Discovery Framework Based on Conditional Generative Model

### v2

@@ Abstract
In this work, we introduce a polymer discovery platform designed to identify polymers with tailored properties efficiently, exemplified through the discovery of high-performance polymer electrolytes. The platform integrates three core components: a conditioned generative model, validation modules, and a feedback mechanism, creating a self-improving system for material innovation. To demonstrate the efficacy of this platform, it is used to identify polymer electrolyte materials with high ionic conductivity. A simple conditional generative model, based on the minGPT architecture, can effectively generate candidate polymers that exhibit a mean ionic conductivity that is significantly greater than those in the original training set. This approach, coupled with molecular dynamics simulations for validation and a specifically designed acquisition mechanism, allows the platform to refine its output iteratively. Notably, after the first iteration, we observed an increase in both the mean and the lower bound of the ionic conductivity of the new polymer candidates. The platform’s effectiveness is underscored by the identification of 19 polymer repeating units, each displaying a computed ionic conductivity surpassing that of Polyethylene Oxide (PEO). The discovery of these polymers validates the platform’s efficacy in identifying potential polymer materials. Acknowledging current limitations, future work will focus on enhancing modeling techniques, validation processes, and acquisition strategies, aiming for broader applicability in polymer science and machine learning.
@@ 1 Introduction
The domain of polymer science is pivotal in shaping advancements across diverse technological arenas. Polymers, with their extraordinary adaptability and customizability, cater to a wide range of applications, spanning from biodegradable materials and high-performance aerospace composites to conducting elements in electronic devices and smart materials in sensor technologies. Notably, polymer electrolytes also play a crucial role, particularly in the field of energy storage [1, 2, 3, 4], exemplifying the versatility of polymers. These applications highlight the immense potential of polymers to be engineered for specific functionalities, such as biocompatibility in medical devices or environmental resilience in various coatings.
@@
Within this realm, conditioned generative modeling presents a particularly relevant technique. By training models on specific conditions or properties, it becomes possible to generate content that meets predetermined criteria. In the current landscape, while conditioned generative models specifically for polymers are still emerging, the concept of integrating machine learning with material science to tailor polymer properties is gaining traction. Our work contributes to this field by introducing a comprehensive polymer discovery platform that leverages the principles of conditioned generative modeling. This platform is not limited to merely suggesting potential polymer candidates but is designed to iteratively improve and refine its suggestions based on continuous feedback and validation. Such a self-improving system embodies a significant leap from traditional methods, offering a more holistic and efficient pathway to polymer material innovation [5, 6, 7].
Specifically, this paper demonstrates the application of our discovery platform in the realm of polymer electrolytes for energy storage technologies. We focus on polymer electrolytes due to their crucial role in the efficiency and safety of devices such as lithium-ion batteries and supercapacitors. Our platform successfully identified polymer electrolytes with ion conductivities superior to the current benchmark, Polyethylene Oxide (PEO), as validated by molecular dynamics (MD) simulations. This achievement not only highlights the potential of our platform in accelerating the discovery of high-performance polymer electrolytes but also sets a precedent for its application across various polymer-based technologies.
@@ 2 Platform
Here, we present a platform structured around an iterative, self-sustaining, and self-improving workflow, comprising three essential components: a conditional generative model, a validation module, and a feedback mechanism. This integrated system allows for continuous refinement and evolution of the discovery process, which we term a "discovery campaign."
The conditioned generative model, the heart of this platform, is tasked with proposing potential polymer candidates. Tailored to incorporate specific target properties, this module is responsible for generating polymers, either by constructing repeating units, oligomers, polymer chains, or 3D structures. For this proof of concept, our focus is on generating the 2D representation of repeat units of polymers. Several aspects heavily influence the performance of the generative model: the (seed) data, the model architecture and hyperparameters, and the training strategies. Different from traditional regression models where a numerical loss is clearly defined, the generative tasks are more ambiguous to evaluate and often require domain knowledge. The process of formulating a comprehensive and domain-relevant evaluation schema and performing the benchmarking across a series of model architectures and training strategies presents its own set of challenges. Our separate research paper [8] addresses these challenges, offering a systematic approach for benchmarking different generative models in different tasks and proposing a comprehensive set of evaluation metrics for the task of generating novel polymers.
Once a batch of polymers is proposed by the generative model, the validation module takes over. This component is responsible for assessing the target properties of the proposed polymers, employing both simulation and experimental validations. For theoretical validation, in the current study, we rely on molecular dynamics (MD) simulations (see Experiment Setup for details). Experimental validation serves as the definitive confirmation for our newly discovered polymers. While this crucial phase falls beyond the scope of the current study, we intend to conduct experimental tests on these candidates in a subsequent phase of our research.
Establishing a feedback mechanism is pivotal to allowing active learning and continuous self-improvement of the model. At the end of each campaign iteration, all validated results are recorded in a database, and strategically sampled for enriching the training data (please see the details of sampling in 3). The model is then retrained to the new data to become increasingly adept at targeting the desired polymers. For future improvements, the uncertainty of the generative model ought to be quantified, and sophisticated acquisition strategies to balance exploration and exploitation are to be tested and implemented.
@@ 3.1 Scientific task
To demonstrate the usage of the platform, we deployed it to identify polymer electrolytes with high ionic conductivity, as advancements in energy storage technologies are increasingly contingent on the development of such materials. Despite significant progress, finding materials that surpass the performance benchmarks set by current standards, such as Polyethylene Oxide (PEO), remains a challenge. PEO, while widely used, falls short of meeting the growing demands for efficiency and stability under varied operational conditions. This limitation is particularly pronounced in applications requiring high charge-discharge cycles and under extreme temperature variations. Our endeavor aims to address this gap by identifying polymer electrolytes with superior ionic conductivities. This quest is not only about surpassing PEO’s performance but also about unlocking new possibilities for safer, more durable, and higher-capacity energy storage solutions.
@@ 3.2 Conditional generation
The core of this demonstration revolves around a conditional generative model based on the minGPT [9] architecture. When trained, this model architecture is capable of completing strings given prompts, just like the popular large language models. The specific model used utilizes the gpt-nano architecture with approximately 120,000 trainable parameters. It is trained on and generates the Simplified Molecular Input Line Entry System (SMILES) [10] code of polymers’ repeating units. Although relatively simple in its design, it is powerful in its application. To direct the model towards polymers with high ionic conductivity, we implemented a method to incorporate the properties of the input during training. This involved the modification of tokenized SMILES strings of known polymer electrolytes, prefixed by their ionic conductivity classes. Specifically, given the range (0.007-0.506 mS/cm) and distribution (mean= 0.062 mS/c, std= 0.036 mS/c) of ionic conductivity in our dataset, we assigned different class labels to high-conductive (top 5%) and low-conductive (lower 95%) polymers and used this class as the leading digit in the input to the model. For example, a polymer with an ionic conductivity of 0.18 (mS/cm) is labeled with class 1 (high conductivity), while another polymer with an ionic conductivity of 0.06 (mS/cm) is labeled with 0 (low conductivity). As the data set evolves over time, the allocation of polymers to the respective high- and low-conductive groups will change accordingly. Additionally, to maintain the importance of the property class in comparison to the lengthy SMILES, and to ensure the model can effectively guide us toward desirable structures, we replicate the property class five times, converting the desired class tokens to "11111". Inspired by the simple design of PEO with a very short repeat unit (OCC) and high ionic conductivity (1.15 mS/cm, at 353K, $Li^{+}.TFSI^{-}$ molality= 1.5 mol/kg), during the iterative polymer generation loops, the model is also biased towards generating small repeating units with SMILES strings containing 10 or fewer tokens. This approach, albeit unrefined, proved crucial for the model’s ability to generate high-conductivity polymers. The reason for the effectiveness of conditioning on short repeat units can be the short distance between negatively charged atoms, such as oxygen atoms, in the polymer backbone that coordinates with Li ions. An effective coordination environment can help with salt dissociation to individual ions and the easier hopping of cations from one coordination site to another. [11, 12]
@@ 3.3 Data
The dataset used in this study is a subset of polymers from High-Throughput Polymer Design - Molecular Dynamics (HTP-MD) database [13, 14], consisting of 6024 linear chain homopolymers. Selected polymers were all unique and composed of H, C, F, S, P, O, N, elements, previously filtered from 53362 structures in the Zinc database [15] to ensure both synthesizability and potential application as electrolytes [16]. The ionic conductivity values for ${}^{+}$${}^{-}$ systems computed from MD simulations performed in the large atomic molecular massively parallel simulator (LAMMPS) [17] with the interaction parameters from Polymer Consistent Force Field (${}^{+}$[18, 19]. The simulations carried out for dataset generation have been previously performed in another study[14], at 353 K and the salt molality of 1.5 mol/kg. More details on the details of simulations and calculation of ionic conductivity can be found in the original work [14, 20].
@@
To skew the model towards high-conductivity polymers, we randomly oversampled the top 5% of polymers in terms of ionic conductivity. This provides a train set that includes the same number of polymers from low-conductive and high-conductive classes. Additionally, PEO was added to the train set and oversampled 4000 times in the seed data. This method of selective oversampling was shown to be instrumental in guiding the model towards generating more promising polymer candidates.
3.4 Validation and feedback
@@
For validation purposes, each iteration of the discovery process involved the generation of 50 polymer candidates, with their ionic conductivities evaluated through molecular dynamics simulations. These simulations adhered to the same protocol used in creating a previous dataset (HTP-MD: [13]). Details of MD simulations, dataset composition, and computing ionic conductivity have been included in  3.3 section, as well as previous studies [16, 14, 20, 13]. To ensure robustness, each candidate underwent five independent simulation replicas to determine its conductivity. Given the randomness in MD simulation results originating from different conformation sampling, this rigorous validation step was crucial for ascertaining the potential of each proposed polymer.
@@
The feedback mechanism of our platform plays a vital role in its iterative learning process. After validation, we add both PEO and newly discovered polymers showing conductivity higher than PEO to the train set and oversample 4000 in total from all newly added polymers. This is to ensure the model can still explore polymers different than PEO. This enriched dataset is then used for retraining the generative model, thereby enhancing its ability to propose increasingly relevant and high-performance polymers in subsequent iterations. For demonstration purpose, the discovery campaign is conducted for two iterations.

### v3

@@ Abstract
In this work, we introduce a polymer discovery platform to efficiently design polymers with tailored properties, exemplified by the discovery of high-performance polymer electrolytes. The platform integrates three core components: a conditioned generative model, a computational evaluation module, and a feedback mechanism, creating a self-improving system for material innovation. To demonstrate the efficacy of this platform, it is used to design polymer electrolyte materials with high ionic conductivity. A simple conditional generative model, based on the minGPT architecture, can effectively generate candidate polymers that exhibit a mean ionic conductivity that is significantly greater than those in the original training set. This approach, coupled with molecular dynamics simulations (MD) for testing and a specifically planned acquisition mechanism, allows the platform to refine its output iteratively. Notably, after the first iteration, we observed an increase in both the mean and the lower bound of the ionic conductivity of the new polymer candidates. The platform’s effectiveness is underscored by the identification of 14 polymer repeating units, each displaying a computed ionic conductivity surpassing that of Polyethylene Oxide (PEO). The performance of these polymers in MD simulations verifies the platform’s efficacy in generating potential polymer candidate materials. Acknowledging current limitations, future work will focus on enhancing modeling techniques, evaluation processes, and acquisition strategies, aiming for broader applicability in polymer science and machine learning.
@@ 1 Introduction
The domain of polymer science is pivotal in shaping advancements across diverse technological arenas. Polymers, with their extraordinary adaptability and customizability, cater to a wide range of applications, spanning from biodegradable materials and high-performance aerospace composites to conducting elements in electronic devices and smart materials in sensor technologies. Notably, polymer electrolytes play a crucial role in the field of energy storage, exemplifying the versatility of polymers [1, 2, 3, 4]. Currently, several challenges are associated with liquid electrolytes, which are the commercially used materials in Li-ion batteries, including flammability [5, 6], toxicity [7, 8], and instability of the electrode-electrolyte interface due to lithium dendrite formation [9, 10]. These issues, along with the several advantages of polymer electrolytes, such as better adaptability to current manufacturing processes compared to ceramics [11] and lower cost compared to ionic liquids [12], make polymer electrolytes promising candidates for future generations of batteries and energy storage devices. These diverse benefits highlight the transformative potential of polymer electrolytes in revolutionizing energy storage technologies, aligning with broader advancements in polymer science for sustainable and high-performance applications.
Recent advancements in the ionic conductivity of polymer electrolytes, crucial for energy storage systems, have been driven by innovative strategies in polymer structure manipulation. These include crosslinking, blending with additives [13, 14], and copolymerization [15, 16, 17]. For instance, He et al. [18] enhanced Li-ion mobility and electrochemical stability by altering the structure of solid polymer electrolytes, moving the carbonate group to the side chain and using a hydrocarbon backbone. This change stabilized the polycarbonate backbone and achieved a conductivity of about 1.1 mS/cm at room temperature with a plasticizer[18]. Similarly, Zhang et al.[19] improved the amorphous domain of polyethylene oxide (PEO) through crosslinking with tetraglyme (TEGDME) and the integration of a rigid linear oligomer, tetraethylene glycol dimethacrylate (TEGDMA), resulting in an electrolyte system that demonstrated an ionic conductivity of $2.7\times 10^{-4}$/cm at room temperature. This system also featured a favorable transference number and low interfacial resistance, showing high potential for energy storage applications. In another approach, Sun et al.[20] developed lithium polysulfides-grafted PEO-based electrolytes via in-situ reduction of sulfur copolymers with poly(ethylene glycol) methyl ether methacrylate (PEGMA). This technique enabled fast Li+ transport and a close electrode-electrolyte interface, achieving dendrite-free Li-metal deposition and maintaining cell capacity over 1200 cycles at 1 C with only a 0.024% capacity decay per cycle. Additionally, Lin et al.[21] designed novel block copolymer electrolytes with three-dimensional networks, enhancing both ionic conductivity to $5.7\times 10^{-4}$/cm and lithium ion transference number to approximately 0.49. They also expanded the electrochemical window to about 4.65V (vs. Li+/Li) and improved mechanical strength at 55 °C by optimizing the ratio of functional units in the copolymers. Despite these significant advances, achieving ion transport properties in solid polymer electrolytes that match those of liquid counterparts remains a formidable challenge, underscoring the ongoing need for innovation in this field.
@@
Within this realm, conditioned generative modeling presents a particularly relevant technique. By training models on specific conditions or properties, it becomes possible to generate content that meets predetermined criteria. In the current landscape, while conditioned generative models specifically for polymers are still emerging, the concept of integrating machine learning with material science to tailor polymer properties is gaining traction. Our work contributes to this field by introducing a comprehensive polymer discovery platform that leverages the principles of conditioned generative modeling. This platform is not limited to merely suggesting potential polymer candidates but is designed to iteratively improve and refine its suggestions based on continuous feedback and evaluation. Such a self-improving system embodies a significant leap from traditional methods, offering a more holistic and efficient pathway to polymer material innovation [22, 23, 24].
In this paper demonstrates the application of our discovery platform in the realm of polymer electrolytes for energy storage technologies. We focus on polymer electrolytes due to their crucial role in the efficiency and safety of devices such as lithium-ion batteries and supercapacitors. We limit the scope of this study to dry (solid) linear chain homopolymers to utilize the HTP-MD [25], a recently developed large database of polymer electrolyte properties computed from MD simulations. While we explored the generation process and compared the performance of different model architectures for the inverse design of polymer electrolytes in our concurrent study [26], this study focuses primarily on integrating the necessary components to propose a discovery framework and evaluate its outcomes. This proof-of-concept study aims to demonstrate that, starting from a set of polymers with computed properties via MD simulations, our framework can design polymers that compete with Polyethylene Oxide (PEO), one of the best-known polymers in dry form in terms of ionic conductivity. We show that our platform successfully designs polymer electrolytes with ion conductivities superior to the current benchmark, PEO, as assessed by molecular dynamics (MD) simulations.
While many other requirements, such as synthesizability, thermochemical stability, compatibility with other components, and in-device performance of newly generated polymer electrolytes, should be considered in the future, the outcome of this study is a small step toward efficiently navigating the large design space. We believe the results of the current study highlight the potential of our platform not only in accelerating the discovery of high-performance polymer electrolytes but also in setting a precedent for its application across various polymer-based technologies.
@@ 2 Platform
Here, we present a platform structured around an iterative, self-sustaining, and self-improving workflow, comprising three essential components: a conditional generative model, a computational evaluation module, and a feedback mechanism. This integrated system allows for continuous refinement and evolution of the discovery process, which we term a "discovery campaign."
The conditioned generative model, the heart of this platform, is tasked with proposing potential polymer candidates. Tailored to incorporate specific target properties, this module is responsible for generating polymers, either by constructing repeating units, oligomers, polymer chains, or 3D structures. For this proof of concept, our focus is on generating the 2D representation of repeat units of polymers. Several aspects heavily influence the performance of the generative model: the (seed) data, the model architecture and hyperparameters, and the training strategies. Different from traditional regression models where a numerical loss is clearly defined, the generative tasks are more ambiguous to evaluate and often require domain knowledge. The process of formulating a comprehensive and domain-relevant evaluation schema and performing the benchmarking across a series of model architectures and training strategies presents its own set of challenges. Our separate research paper [26] addresses these challenges, offering a systematic approach for benchmarking different generative models in different tasks and proposing a comprehensive set of evaluation metrics for the task of generating alternative polymers.
Once a batch of polymers is proposed by the generative model, the evaluation module takes over. This component is responsible for assessing the target properties of the proposed polymers, employing both simulation and experimental validations. For computational evaluation, in the current study, we rely on molecular dynamics (MD) simulations (see Experiment Setup for details). Experimental validation serves as the definitive confirmation for our newly discovered polymers. While this crucial phase falls beyond the scope of the current study, we intend to conduct experimental tests on these candidates in a subsequent phase of our research.
Establishing a feedback mechanism is pivotal to allowing active learning and continuous self-improvement of the model. At the end of each campaign iteration, all computed results are recorded in a database, and strategically sampled for enriching the training data (please see the details of sampling in 3). The model is then retrained to the new data to become increasingly adept at targeting the desired polymers. For future improvements, the uncertainty of the generative model ought to be quantified, and sophisticated acquisition strategies to balance exploration and exploitation are to be tested and implemented.
@@ 3.1 Scientific task
To demonstrate the usage of the platform, we deployed it to identify polymer electrolytes with high ionic conductivity, as advancements in energy storage technologies are increasingly contingent on the development of such materials. Despite significant progress, finding materials that surpass the performance benchmarks set by current standards, such as Polyethylene Oxide (PEO), remains a challenge. PEO, while widely used, falls short of meeting the growing demands for efficiency and stability under varied operational conditions. This limitation is particularly pronounced in applications requiring high charge-discharge cycles and under extreme temperature variations. Our endeavor in the current study mainly aims to demonstrate the power of this discovery framework for identifying polymer electrolytes with superior ionic conductivities. While this study does not specifically target thermal stability or varied operational conditions, the framework has the potential to be extended in future work to address these critical challenges, potentially leading to safer, more durable, and higher-capacity energy storage solutions.
@@ 3.2 Conditional generation
The core of this demonstration revolves around a conditional generative model based on the minGPT [27] architecture. When trained, this generative model is capable of generating a polymer electrolyte candidate given a lead prompt signaling what kind of property is needed (with high or low ionic conductivity), just like the popular large language models. The specific model used utilizes the gpt-nano architecture with approximately 120,000 trainable parameters, and we used the Simplified Molecular Input Line Entry System (SMILES) [28] code of polymers’ repeating units to represent the polymer electrolyte candidates. Although relatively simple in its design, it is powerful in its application. To direct the model towards polymers with high ionic conductivity, we implemented a method to incorporate the properties of the input during training. This involved the modification of tokenized SMILES strings of known polymer electrolytes prefixed by their ionic conductivity classes. Specifically, given the range (0.007-0.506 mS/cm) and distribution (mean= 0.062 mS/cm, std= 0.036 mS/cm) of ionic conductivity in our dataset, we assigned different class labels to high-conductive (top 5%) and low-conductive (lower 95%) polymers and used this class as the leading digit in the input to the model. Therefore, the dividing line between the "high-conductivity" and "low-conductivity" categories is set to be 0.012 mS/cm, and this is fixed throughout the experiment. For example, a polymer with an ionic conductivity of 0.18 (mS/cm) is labeled with class 1 (high conductivity), while another polymer with an ionic conductivity of 0.06 (mS/cm) is labeled with 0 (low conductivity). As the data set evolves over time, the allocation of polymers to the respective high- and low-conductive groups will change accordingly. Additionally, to maintain the importance of the property class in comparison to the lengthy SMILES, and to ensure the model can be effectively guided towards desirable structures, we replicate the property class five times, converting the desired class tokens to "11111", which is shown to result in better polymer electrolyte candidate empirically. Therefore, the effective prompt that signals the conditioned generative model to generate polymer electrolyte candidates with high ionic conductivity is simply a leading string of "11111", and the model will then complete the string by generating a SMILES code that represents a specific polymer candidate.
Inspired by the simple design of PEO with a very short repeat unit (OCC) and high ionic conductivity (1.15 mS/cm, at 353 K, $Li^{+}.TFSI^{-}$ molality= 1.5 mol/kg), during the iterative polymer generation loops, the model is also biased towards generating small repeating units with SMILES strings containing 10 or fewer tokens. This approach, albeit unrefined, proved crucial for the model’s ability to generate high-conductivity polymers. The reason for the effectiveness of conditioning on short repeat units can be the short distance between negatively charged atoms, such as oxygen atoms, in the polymer backbone that coordinates with Li ions. An effective coordination environment can help with salt dissociation to individual ions and the easier hopping of cations from one coordination site to another [29, 30]. Regardless of the imposed restriction on the length of repeat units (10 and fewer characters), the model itself shows a tendency toward generating short repeat units, which is mainly due to oversampling PEO that will be discussed in Section 3.4.
3.3 Model details
The generative model we implemented here is based on a minimal implementation of the GPT model. [27]. The model first converts a sequence of tokens (SMILES vocabulary) to two embeddings, including token embedding and positional embedding. Token embedding represents the meaning of individual words or symbols in a high-dimensional space, while positional embedding encodes the order or position of tokens within a sequence to provide contextual information to the model. These embeddings are then passed through multiple layers of transformer blocks. Each transformer block mainly consists of a multi-head masked self-attention layer and a feed-forward neural network, following the original transformer architecture [31]. The loss function is cross entropy loss comparing tokens in the generated and actual sequences.
We performed a grid search to tune the hyperparameters of minGPT and optimize the performance of the model. The mean values of six metrics (chemical validity, uniqueness, novelty, synthesizability, similarity, diversity) are utilized as the evaluation metric for the model’s performance. It is important to note that in this study, novelty is defined as a polymer structure that has not been seen by the model during the training process and is not based on literature novelty. We used 3 independent hyperparameters, including "training epoch," "model architecture," and "temperature." The "training epoch" varies from 1000 to 10000 with a uniform interval of 1000. The "model architecture" corresponds to HuggingFace’s Transformer-based model architecture. [32] We select 3 different architectures which are "gpt-2", "gpt-mini" and "gpt-nano". In terms of the "temperature", the values are from 0.1, 0.5 and 1.0. A higher temperature results in higher "creativity" for the GPT model. For the optimal performance, the "training epoch" is 6000, the "model architecture" is "gpt-nano" and the "temperature" is 1.0. As a result, the number of attention heads in the minGPT model is 3, and the hidden dimensions are set to 48, which results in around 0.12M parameters. The hyperparameter selection and the grid search results for the GPT model can be in the Supporting Information Table. S1, S2, S3, and S4. For more details, please refer to our concurrent work [26].
@@ 3.4 Data
The dataset used in this study is a subset of polymers from High-Throughput Polymer Design - Molecular Dynamics (HTP-MD) database [33, 25], consisting of 6024 linear chain homopolymers. Selected polymers were all unique and composed of H, C, F, S, P, O, N, elements, previously filtered from 53362 structures in the Zinc database [34] to ensure both synthesizability and potential application as electrolytes [35]. The ionic conductivity values for polymer-Li+.TFSI- systems computed from MD simulations performed in the large atomic molecular massively parallel simulator (LAMMPS) [36] with the interaction parameters from Polymer Consistent Force Field (PCFF+[37, 38]. The PCFF+ force field has been employed in previous studies to explore various properties of electrolyte systems [35, 39, 40, 41, 42, 43, 44]. Additionally, multiple studies have compared the molecular dynamics (MD) simulation results obtained using the PCFF+ force field with experimental data and density functional theory (DFT) calculations for both polymer [35, 45, 46] and liquid [47, 48, 49] electrolytes. The simulations carried out for dataset generation have been previously performed in another study [25], at 353 K and the salt molality of 1.5 mol/kg. More details on the details of simulations and calculation of ionic conductivity can be found in the original work [25, 50].
@@
To skew the model towards high-conductivity polymers, we randomly oversampled the top 5% of polymers in terms of ionic conductivity. This provides a train set that includes the same number of polymers from low-conductive and high-conductive classes. Additionally, PEO was added to the train set and oversampled 4000 times in the seed data. This method of selective oversampling was shown to be instrumental in guiding the model towards generating more promising polymer candidates. It should be noted that in the current study, oversampling was performed by including duplicate SMILES strings, and we have not tried other methods, such as randomization [51].
3.5 Evaluation and feedback
@@
For evaluation purposes, each iteration of the discovery process involved the generation of 50 polymer candidates, with their ionic conductivities evaluated through molecular dynamics simulations. These simulations adhered to the same protocol used in creating a previous dataset (HTP-MD: [33]).
We carried out molecular dynamics (MD) simulations on polymer-(Li+.TFSI-) systems at a temperature of 353 K and a salt concentration of 1.5 mol/kg using LAMMPS [36] with the Polymer Consistent Forcefield (PCFF+) [37]. The simulation process included initial relaxation and equilibration of the polymer-salt mixture, followed by a production phase to gather data for computing ion transport properties, such as ionic conductivity. The equilibration stage involved running sequential NVT (constant number of particles, volume, and temperature) and NPT (constant number of particles, pressure, and temperature) ensembles to achieve densities close to theoretical values. For the production run, an NVT simulation at 353 K was conducted for 5 ns with a 2.0 fs time step. The resulting trajectories were then analyzed to compute ion transport properties using the cluster Nernst-Einstein equation [39]. The analysis code used to compute ionic conductivity is consistent with the one used to generate the HTP-MD database and is available at https://github.com/TRI-AMDD/htp__\__md. More details about MD simulations, dataset composition, and computing ionic conductivity have been included in  3.4 section, as well as previous studies [35, 25, 50, 33]. The above-described simulation protocol has been previously employed to assess the computed ion transport properties of polymers in the HTP-MD database and compare them with the experimentally measured ionic conductivity values. [35]
The repeat units of the generated polymers are polymerized to have at least 150 heavy atoms (non-H) in their backbone, with the two ends terminated by methyl groups. This approach is consistent with the method used to generate the HTP-MD database, allowing us to compare the performance of newly generated polymers with the training set. Also, to ensure robustness, each candidate underwent five independent simulation replicas to determine its conductivity. Given the randomness in MD simulation results originating from different conformation sampling, this rigorous step was crucial for ascertaining the potential of each proposed polymer.
@@
The feedback mechanism of our platform plays a vital role in its iterative learning process. After evaluation, we added both PEO and newly discovered polymers showing conductivity higher than PEO to the train set and oversampled 4000 in total from all newly added polymers. This is to ensure the model can still explore polymers that are different from PEO. This enriched dataset was then used to retrain the generative model, thereby enhancing its ability to propose increasingly relevant and high-performance polymers in subsequent iterations. For demonstration purpose, the discovery campaign is conducted for two iterations.